# Collective steering: Tracer-informed dynamics

### v1

@@ 6 Concluding remarks
The theme of the present work brings up a hitherto unstudied angle in optimal transport problems, where, now, additional data reveal the dynamics of internal degrees of freedom in the flow. Paradigmatically, in either fluid flow or the flow of a flock of birds, tagging specific individuals (tracer particles or birds, respectively) may provide information about rotational components of the collective dynamics that are not reflected in the aggregate distributions.
The work herein explores this theme in the elementary setting of linear dynamics and Gaussian distributions. From a controls perspective, we envision stochastic formulations, where process and measurement noise corrupt the available data, that mirror the deterministic setting presented herein. From an applications standpoint, tracer information is naturally available in several fields. For instance, in computer vision, tracers may correspond to landmarks or facial features, and, in fluid flows, tracers are often injected for precisely the same reasons as postulated herein. In these and other applications, we expect that the conceptualization present in this paper will prove relevant.

### v2

@@ 5 Concluding remarks
The theme of the present work brings up a hitherto unstudied angle in optimal transport problems, where additional data now reveal the dynamics of internal degrees of freedom in the flow. Paradigmatically, in either fluid flows or the flows of a flock of birds, tagging specific individuals (tracer particles or birds, respectively) may provide information about rotational components of the collective dynamics that are not reflected in the aggregate distributions.
This work explores the theme in an elementary setting of linear dynamics and Gaussian distributions. From a control perspective, we envision stochastic formulations, where measurement noise corrupts the available data, that mirror the deterministic setting presented herein. From an applications standpoint, tracer information is naturally available in several fields. For instance, in computer vision, tracers may correspond to landmarks or facial features, and, in fluid flows, tracers are often injected for precisely the same reasons as postulated herein. Moreover, it is of interest to explore how the conceptualization presented can be utilized in related developments, such as [25, 26, 27, 28].